x + y = -2
2x - y = 5

Answer by stanbon(75887)

x + y = -2
2x - y = 5
--------------
Add the two equations and solve for "y":
3x = 3
x = 1
------
Since x+y = -2, 1 + y = -2 ; y = -3
=====================================
Ans: (1,-3)
========================
Cheers,
Stan H.